UT 20131112 LUCI2 Nighttime Commissioning

Observer: WSeifert, JHeidt, PBuschkamp, APramskiy, MEdwars, DThompson,KPolsterer
Telescope Operator: DGHuerta
LBTO Support: DMiller, TSargent

Plan for the night

Weather is photometric, we will take advantage of that

  • Twilight flats
  • Check update of SW (wait for collimation before data taking starts)
  • Guide probe transformation verification (if seeing is good by DThompson, DMiller)
  • Telescope focus using TT (with DMiller)
  • Photometric ZP's in all filters
  • Extinction for BB-filters (JHKKs across a range of AMs)
  • Photometric gradient across detector (one BB and one nB-filter, KS, Br_gam)
  • PR-type object, test for large offsets
  • Test of LS-acq using Ingo's new tool
  • Twilight flats in the morning

Summary

Photometric night: spend most of the night on photometric ZP's as function of AM

New SW-version, data taking now starts once telescope is collimated

AGw transformation verification failed

Tried LS-acquisition but had problems with the SW => postponed

Preparations

DThompson changed the AGw2-focus by -1.376mm, GCS restartet
=> should be good in focus

MEdwards did a great job in creating scripts all night

Details

Files are luci2.20131111.iiii

00:30 Open dome

00:31 Sunset, weather photometric, low wind, temp +5deg

00:32 New SW-patch installed

00:40 Ks-flats #1-18 20000cts, no real flats - too late -> ReadoutManager is not updating, restart doesn't help

01:13 David sets up the telescope

Photometric standards

01:41 Preset BS9188, AM=1.07, JHKKS, preset failed, moon too close

01:55 Preset to BS9101, AM=1.326, JHKKs, #23-27, stopped unexpectedly after 5th exp.
when it should move to H-filter

02:10 Seeing variable and not very good. Ranges from 0.7-1.5" on the guider

02:14 Continue on BS9188 in HKKs-filter, #28-42

02:38 BS9182, AM=1.5, JHKKs, new set of scripts, do now 2*5sec integrated, 5 pos per filter,
problems in collimating, moon too close (as for most Pearsson-stds vailable at the moment)

02:55 Redo BS9182, David has collimated on-axis beforehand, nope, doesn't help

02:59 BS9105, AM=1.7, JHKKs, #46-65 , FWHM=0.8-1"
The first script having 20 positions and 4 filters who went through

03:25 BS9188 again, geometry between Moon and BS9188 has changed, see if it works now,
David collimates on-axis before, than stop for new SW-version

03:58 New SW-version in place, includes now "wait for collimation" before start of data taking
and predefined "save/frame type" in readout manager, presert BS9188, JHKKs, AM=1.01,
"wait for collimation successfull, but error message for instrument manager* => go back to old version

04:12 BS9188 with old SW-version, JHKKs, AM=1.01, problem with instrument manager, stopped

04:23 Deployed new SW-version again (elements need to be initialized), BS9188, seems to work, files #71-90,
JHKKS, AM=1.03, FWHM=0.6", seeing seems to have stabilized, Duty Cycle 30%
Script went through using the new SW-version

04:36 BS9187, JHKKs, AM=1.6, does not collimate, LUT for secondary not sufficiently good
at low elevations, good done with help by DMiller, started at AM=1.7, #91-110, images elongated

AGw transformation verification

05:20 Dave and Doug took over to start verification of the new AGw transform

Stone_B1 field
0111 test w/BrGamma 2x5.0s
0112 start with star 2
  • could not find third star
  • pointing check...OK
  • restart transform_collect
0113 star 2
0114 star 4 (skipped star 3)
0115 star 5
0116 star 6
0117 star 7
0118 star 8
0119 star 9
Aborted transform check...on-axis source not in a constant position.
DMiller and DThompson discussed problem, will need to re-run the creation of the transform again with different parameters.
Problem was the corrections derived from the pointing mask were applied in the wrong direction.

ZP's for all filters

06:20 BS9107, script BS9107-FW2, JHKKsz, AM=1.03, #120-144, stars partly elongated
Background: Jband: 1000cts/sec, K-band 1600cts/sec, Ks 800cts/sec, z=60cts/sec

06:41 BS9107, script BS9107-FW1, stopped at Y2

07:07 BS9107, reconvened at Y2, #145-204

07:38 BS9107, zJ_spec, HK_spec, #205-206, script stopped after 2nd exposure

07:54 BS9107, restart, zJ_spec, HK_spec, #207-216

08:04 BS9188, JHKKs, preset fails, recheck pointing, OK; #217-236, FWHM ~ 0.7" @ 30deg elevation

08:25 BS9105, JHKKs, AM=1.3, #237-256

Photometric gradient across detector

08:50 BS9107, Br_gam/K, positions across detector, preset failed, check pointing,#257-275

LS-acquisition

09:28 BS9107, script BS9107_longslitACQ_PA0, first attempt to use IStilz new LS acquisition tool,
does first sky and than target-image followed by through-slit ima 1"-slit,
....lets see how it goes, script got stuck (problem with instrument manager)
=> inconsistency of with naming of mask scriptor/MOS-config, reconfigured, redo
=> MOS-panic when moving 1.0" LS to turnout, reason: see log-message 108871, rot=-83, ele=56

10:10 Changed to LS1.5_900 & redo, got stuck again

10:46 Another attempt, got again stuck, this time with instrument manager, Kai & Alexander working on fix

12:30 Problems with SW not fixed yet, in the meantime seeing considerably degraded (>2"),
=> Done for the rest of the night (except flats)

13:16 J-flats: #281-312, 4000-25000cts
13:23 H-flats: #313-374, 4500-22000cts
13:29 Ks-flats: #375-504, 8300-21000cts

13:39 Close dome

-- JochenHeidt - 12 Nov 2013
Topic revision: r10 - 05 Dec 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback